Uplinked the daily Master Command Loads and Ephemeris.
As part of the spin-up to the 40 rpm operational rate, the AMSR-E sensor rotation was increased in two phases: from 4 to 12 rpm (1440z), then from 12 to 18 rpm (1750z).
The USO-1 clock was adjusted to the slowest frequency, and USO-2 clock drift rate was adjusted. The clocks will be monitored over the next 3-4 day period as the adjustments take effect.
As part of planned check out activities, the AMSU-A1 and A2 were transitioned from their Full Scan mode to their Parked mode.
The HSB scanner was transitioned from Scan Normal mode to Park as part of their planned checkout activities.
Science data processing status: HSB Level 0 science data has been received by the JPL DAAC. Processing of the HSB data to Level 1A has been delayed until after the initial AMSU data processing is complete.
The following activities are scheduled for the next 24 hours:
Activities will be performed to collect data for analysis and confirmation of suitable conditions prior to continuing the AMSR-E spin-up to 40 rpm.
Aqua Spacecraft Status:
Spacecraft State is currently Standby.
GNC Mode is Fine Point Mode, HiFi Ephemeris.
Aqua Instrument Status:
AIRS is in Decontaminate state, Quiet bus is ON, Noisy bus is ON, Cryocooler is OFF, Scan mirror rotating is OFF and earth shield is CLOSED.
AMSR-E is in Run-Up mode, rotating at 21 RPM.
AMSU A1 is in Parked mode and the antenna is in the Space View position.
AMSU A2 is in Parked mode and the antenna is in the Space View position.
CERES Aft is in Safe mode and the covers are CLOSED.
CERES Fore is in Safe mode and the covers are CLOSED.
HSB is in Park mode and the antenna is in the Space (pos1) position.
MODIS is in Standby mode, Nadir Door is Unlatched and CLOSED, Solar Diffuser Door is Unlatched and CLOSED, and the Space View Door is Unlatched and CLOSED.
The apparent misalignment of approximately 0.2 degrees between Star Tracker Assembly (STA) 1 and STA2 has been resolved. Ball has confirmed that bore-sight rotation adjustment sign convention was not correct. This sign reversal accounts for the shif relative to the expected STA1 to STA2 alignment. The matrices will be revised and reloaded this coming week.
A patch has been developed to address the timing issue between the CTC and the USO. The patch was successfully loaded and demonstrated 5/16/02 @ 0200Z. USO1 was commanded to its slowest frequency to reduce the large positive GIRD clock error. Clock is currently drifting at 28 ms/day. GIRD clock error expected to be near zero in approximately 3.8 days. At that time, the frequency will be adjusted to produce a small negative clock drift. USO2 frequency has been adjusted resulting in a GIIS clock drift of 0.5 ms/day. In addition, the GIIS clock was time jammed resulting in a clock error of 1.5 ms. The GIIS clock can now be maintained per standard operating procedures to within + or 10 ms.
During the AMSR-E run up from 4 to 12rpm and 12 to 18rpm, GNC detected disturbances in roll and yaw rates. The yaw amplitude remains constant from the 12 rpm run up which is consistent with a static imbalance. The increasing peak to peak amplitude in roll may be due to a solar array excitation mode. As an attempt to avoid the current SA excitation mode (at 18 rpm), AMSR-E was run up to an intermediate speed of 21 rpm. The AMSR-E OBM positions were re-verified, and a demonstration of OBM movement (OBM-XH) was performed. GNC will be preparing to perform necessary OBM balancing upon sufficient data gathering/analysis.
AMSR-E was unable to reset their SPC error flag 1 and SPS error flag 4. Resolution is under investigation.